# Large Language Model LRM Large Reasoning Model

### v2

@@ 4.2 Observation Masking: Dominant Efficiency with Minimal Complexity
Having established the clear need for context management, we now turn to the central finding of our work: the surprising power of simplicity. As we can see in Table˜1, in four out of five experimental setups Observation Masking yielded the lowest cost per instance. It achieves this by drastically reducing the number of environment observation tokens processed in each agent turn without incurring the computational overhead of a separate summarization call. This is highly effective, because the agent’s context is heavily skewed towards environment observations in SE (Figure˜1).
Let’s consider our best-performing model Qwen3-Coder 480B Team (2025); Hui et al. (2024). For it, the Observation Masking strategy achieves a cost of \$0.61 per instance, a 52.7% reduction compared to the Raw Agent baseline (\$1.29) and a tangible improvement over LLM-Summary (\$0.64) that amounts to 15\$ in savings across 500 instances in SWE-bench Verified Chowdhury et al. (2024). This demonstrates that even for state-of-the-art models, a simple omission strategy is a powerful baseline for efficient context management.
@@ 4.3 Challenging the Need for Complex Summaries
In fact, Observation Masking not only competes with LLM-Summary, but can outperform it. With the Qwen3-Coder 480B Team (2025); Hui et al. (2024) model, Observation Masking achieved a solve rate of 54.8%, a slight improvement over the LLM-Summary’s 53.8%. Similarly, for Gemini 2.5 Flash Comanici et al. (2025) (thinking), it outperformed LLM-Summary by five percentage points. In the cases where LLM-Summary did perform better, such as with Gemini 2.5 Flash, the margin was minimal (36.0% vs. 35.6%). This indicates that Observation Masking consistently performs on-par with, or better than the LLM-Summary strategy. Moreover, using one of the two tested context management strategies outperformed Raw Agent in three out of five configurations. Highlighting that increased context can eventually result in noise for LLMs Lindenbauer et al. (2025); Hong et al. (2025); Liu et al. (2024).
The implication is clear: the most recent context is often sufficient for code-generating agents. Retaining the entire history, or even a sophisticated summary of it, may not be the most effective use of the model’s limited context window and our research budget.
@@ 4.4 The Trajectory Elongation Effect
Figure 4: Impact of context management strategies on trajectory length. Box plots show the distribution of trajectory lengths (in turns) across different strategies within SWE-agent Yang et al. (2024). LLM-Summary consistently leads to longer trajectories, suggesting they mask failure signals that would otherwise prompt earlier termination. The star indicates the mean trajectory length.
@@ 5 Discussion
Our results consistently demonstrate that a simple Observation Masking strategy is not only more cost-effective but also surprisingly competitive in solve rate compared to more complex LLM-Summary methods within SWE-agent Yang et al. (2024). This challenges the prevailing assumption that sophisticated, semantic context compression is always superior. The discovery of the "trajectory elongation" effect (Section˜4.4) provides a compelling explanation for this phenomenon: the smoothed nature of a summary may inadvertently mask signs of a failing trajectory, encouraging the agent to persist in unproductive loops. We hypothesize the mechanism generalizes across scaffolds that trigger summarization by turn count or context thresholds, but we leave verification to future work.
5.1 Critic-Enhanced Summarization
A natural follow-up question is whether the LLM-Summary strategy could be improved by making the summarization process more intelligent. We explore this by enhancing our LLM-Summary strategy with execution-free feedback, a technique that has shown promise in scaling test-time compute for SE agents Pan et al. (2024); Jain et al. (2025); Zhang et al. (2025); Antoniades et al. (2025). In this approach, the LLM simultaneously generates a summary and critical analysis of the trajectory, incorporating both into the compressed context. This is akin to providing reflections within a single rollout, instead of across multiple rollouts Shinn et al. (2023).
Testing on 150 samples from SWE-bench Verified Chowdhury et al. (2024) using SWE-agent Yang et al. (2024), this critic-enhanced approach using the prompt presented in Figures˜12, 13 and 14 showed no improvement in solve rate over standard LLM-Summary. More concerning, we observed exacerbated trajectory elongation patterns, with critic-enhanced runs producing even longer trajectories than standard summarization. This is perhaps unsurprising: the critic’s reflections naturally encourage the agent to explore alternative solution paths, try additional debugging strategies, or reconsider its approach, all of which translate to more turns, thus driving cost and reducing efficiency gains.
This finding reinforces our central insight about trajectory elongation. While execution-free feedback aims to improve agent decision-making, it paradoxically increases computational costs by extending exploration. The critic’s guidance, rather than helping the agent efficiently recognize dead ends, provides additional avenues to pursue, further delaying termination. Furthermore, this increased cost, does not lead to increased downstream performance. This suggests that effective memory systems for AI agents require fundamental rethinking: simply adding more sophisticated feedback to summaries may compound rather than solve the efficiency challenges we identify. See Section˜D.3.
@@ 5.2 The Costs of Summarization
A closer examination of the cost breakdown reveals that the efficiency gap between strategies stems from two complementary effects. First, the "trajectory elongation" effect discussed in Section˜4.4, and second, the costs of generating the summary.
As shown in Table˜2, the direct API cost of generating summaries accounts for up to 7.2% of the total instance cost. Importantly, these summarization calls are particularly expensive because each requires processing a unique sequence of turns, limiting cache reuse to the LLM-Summary system prompt (Figure˜11). This poses a critical limitation given that, several modern LLM APIs (e.g., Gemini) offer substantially cheaper cache hits than cache misses (often on the order of  3–4× cheaper). Once we subtract these summarization costs from the total, the efficiency difference between LLM-Summary and Observation Masking largely disappears for most experiments. This indicates that the summarization API calls themselves constitute a significant portion of the efficiency gap. Nonetheless, the more complex LLM-Summary strategy is still unable to consistently outperform the remarkably strong Observation Masking strategy on cost-efficiency, highlighting the need for more sophisticated strategies, potentially combining the strengths of these two approaches.
@@ 6 Limitations
While our study provides a rigorous evaluation of context management strategies, its scope has three main limitations.
First, we conduct our experiments within a single scaffold (SWE-agent Yang et al. (2024)). We do not claim generality across agent scaffolds. Preliminary results on SWE-bench Verified-50 Badertdinov et al. (2024); Chowdhury et al. (2024) using OpenHands Wang et al. (2025) indicate scaffold-specific nuance to our findings. See Appendix˜E for details. We leave systematic evaluation of further agents, such as OpenHands Wang et al. (2025), to future work.
Second, our experiments are situated exclusively within the SE domain, using the SWE-bench Jimenez et al. (2024) benchmark. This domain is characterized by long, verbose tool outputs, a condition that naturally favors the efficiency of Observation Masking. Consequently, our findings on the superiority of this strategy may not generalize to domains where agent-environment interactions are more succinct.
Finally, both the LLM-Summary and Observation Masking strategy, are triggered by simple, non-adaptive heuristics. Observation Masking employs a fixed-size rolling window that is agnostic to the relevance or staleness of past observations (e.g., retaining a file’s content after it has been modified). Similarly, LLM-Summary operates on a fixed turn-based schedule, ignoring semantic boundaries or agent subgoals. Our study also does not explore hybrid strategies that could potentially combine the benefits of both approaches.
@@ 7 Conclusion
Our comprehensive study on context management strategies within SWE-agent Yang et al. (2024) establishes the critical need for context management for economically feasible, and environmentally sustainable deployment. Foregoing context management more than doubles execution costs. Yet, contrary to the trend toward increasingly complex summarization, a simple Observation Masking strategy consistently performs on par with a sophisticated LLM-Summary approach in cost efficiency and solve rate within SWE-agent Yang et al. (2024).
Our analysis uncovers a fundamental insight: LLM-based summarization can inadvertently encourage “trajectory elongation,” where agents persist on unproductive paths, an effect especially pronounced with powerful models like Gemini 2.5 Flash or Qwen3-Coder 480B. Moreover, summary-generation LLM calls account for a nontrivial fraction of per-instance cost for the strongest models (5–7%), eroding the net efficiency of these approaches. These observations point to several concrete next steps: evaluate whether these trends generalize to other agent scaffolds (e.g., OpenHands Wang et al. (2025)); move beyond chronological compression toward data-driven selection of what to retain; explore hybrid schemes that primarily use Observation Masking but invoke selective LLM summarization at truly critical junctures; learn triggers that decide when summarization is beneficial versus harmful (e.g., detecting looping or plateaus); evaluate the generality of these strategies in domains beyond SE; and reduce summarization overhead via lightweight specialist summarizers or distillation.
Overall, our work demonstrates that in the quest for more capable AI agents, complexity is not always the answer. Sometimes the simplest solution is also the most effective.
@@ Appendix A Experimental Configuration
To compute the instance cost reported in our experiments, we distinguish between proprietary models we access through the API, and models we host locally on our infrastructure. For the Gemini Comanici et al. (2025) experiments, which we access through the Vertex AI API, we report the cost returned by the API. For Qwen Team (2025); Hui et al. (2024) experiments, we self-host the models as described in Section˜3.2. To compute the cost, we use the observed per-turn token usage and post-hoc compute the cost based on the official Alibaba API pricing666 https://www.alibabacloud.com/help/en/model-studio/models#16ff9753e1ctz. Note that in contrast to the Vertex AI platform, the official Alibaba API pricing for the Qwen3-32B model, does not distinguish between cache hit and miss input tokens. This leads to an inflated cost of the Qwen3-32B experiments. We refer readers interested in the cost structure of our experiments under different pricing schemes to out HuggingFace dataset [^7] Data: https://huggingface.co/datasets/the-complexity-trap-19fna01lf1Y/main-experiments, in which we release all experimental data for our main experiments and our repository [^8] Code: https://anonymous.4open.science/r/complexity-trap-6167/.
Appendix B LLM Summary Prompt
We share our prompt template for summary generation in Figure˜11. Compared to OpenHands Wang et al. (2025), we remove the part of the prompt that aims to handle summary generation for tasks outside the SE domain, since out work is purely focused on the SE domain. In addition to the system prompt shown in Figure˜11, we provide a joint critic-summarization prompt in Figures˜12, 13 and 14. We discuss the effects of generating execution-free in Section˜D.3 and Section˜5.
@@ Appendix C Detailed Main Results
Table 3: Comparison of context management strategies with 95% bootstrap confidence intervals, showing asymmetry. We use † to indicate significance compared to the Raw Agent. We report Solve Rate (effectiveness, $\uparrow$) and Instance Cost (efficiency, $\downarrow$). For each model, we boldface the best-performing context management strategy for each metric (relative to the Raw Agent baseline). Change is reported relative to the Raw Agent baseline. All experiments use SWE-agent Yang et al. (2024) on SWE-bench Verified Chowdhury et al. (2024).
@@ Appendix D Additional Studies
For the critic-enhanced summarizer in Figure˜6, we experiment on SWE-bench Lite-50 Badertdinov et al. (2024). For the sensitivity to the rolling window size $M$ of the Observation Masking strategy and the configurations of the LLM-Summary strategy, we show our results on a randomly sampled 150-instance subset of SWE-bench Verified Chowdhury et al. (2024) that we release with our code. We conduct these studies with GPT-4.1-mini OpenAI et al. (2024).
Figure 5: Downstream task performance of a single experiment on a randomly generated 150-sample subset of SWE-bench Verified Chowdhury et al. (2024) across various different configuration combinations with respect to the tail length $M$. We find that a larger summarization window compared to the tail length improves performance.
D.1 Rolling Window Size M
We experiment with the rolling window size $M$ of the Observation Masking strategy. In Figure˜9 we can see that the performance of the strategy peaks at $M=10$, before falling again when further increasing the window size to $M=20$. Thus, we use this configuration of the Observation Masking strategy for our main experiments.

### v3

@@ 4.2 Observation Masking: Dominant Efficiency with Minimal Complexity
Having established the clear need for context management, we now turn to the second central finding of our work: the surprising power of simplicity. As we can see in Table˜1, in four out of five experimental setups Observation Masking yielded the lowest cost per instance. It achieves this by drastically reducing the number of environment observation tokens processed in each agent turn without incurring the computational overhead of a separate summarization call. This is highly effective, because the agent’s context skews heavily towards environment observations in SE (Figure˜1). Furthermore, this strategy requires fewer warm-up turns than LLM-Summary ($M=10$ vs $N+M=31$). This results in quicker and more robust cost reductions, even on short trajectories (e.g., Qwen3-32B (thinking), see Appendix˜B).
Notably, this finding holds across model configurations. Furthermore, while \$0.03 in cost reductions between LLM-Summary and Observation Masking for Qwen3-Coder 480B seems small, it already amounts to \$15 across the entire benchmark. This highlights that even small cost-efficiency gains can have a significant impact on the economic viability or large-scale LLM agent deployments and underscores the need for research on efficient context management.
@@ 4.3 Challenging the Need for Complex Summaries
In fact, Observation Masking not only competes with LLM-Summary, but can outperform it. With the Qwen3-Coder 480B [27, 8] model, Observation Masking achieved a solve rate of 54.8%, a slight improvement over the LLM-Summary’s 53.8%. Similarly, for Gemini 2.5 Flash [6] (thinking), it outperformed LLM-Summary by five percentage points. In the cases where LLM-Summary did perform better, such as with Gemini 2.5 Flash, the margin was minimal (36.0% vs. 35.6%). This indicates that Observation Masking consistently performs on-par with, or better than the LLM-Summary strategy.
The implication is clear: the most recent context is often sufficient for SE agents. Retaining the entire history, or even a sophisticated summary of it, may not be the most effective use of the model’s limited context window and our research budget.
@@ 4.4 The Trajectory Elongation Effect
Figure 4: Impact of context management strategies on trajectory length. Box plots show the distribution of trajectory lengths (in turns) across different strategies within SWE-agent [35]. LLM-Summary consistently leads to longer trajectories, suggesting they mask failure signals that would otherwise prompt earlier termination. The star indicates the mean trajectory length.
@@ 5 Discussion
In this section we probe the generality of our findings on OpenHands [30], discuss the impact of the LLM summary generation on the cost structure of the LLM-Summary strategy, and introduce a novel hybrid approach combining the strengths of both strategies.
5.1 Probing for Generality With OpenHands
To investigate the generality of our main results across scaffolds, we probe for generality with OpenHands on a 50-instance slice of SWE-bench Verified [5, 3] using Gemini 2.5 Flash [6] (no thinking), with a turn limit of 250, LLM-Summary N=21, M=10, and Observation Masking M=10 and M=58. We present the results Figure˜5a.
First, we note that the Observation Masking rolling window size $M$ is an agent-specific hyperparameter that requires tuning. If we simply re-use the optimal value from SWE-agent [35], Observation Masking performance degrades drastically. However, after tuning, we can reproduce our results on this agent scaffold. We hypothesize that we need to tune this hyperparameter due to scaffold-specific implementation details. For example, SWE-agent [35] directly elides retries due to syntax errors from the dialog history. However, OpenHands [30] retains such retry turns. This means we need a larger window size to retain an informative context for this agent.
Overall, this provides initial evidence that our findings generalize across agent scaffolds if the agent’s context similarly skews toward environment observations, as typically is the case in SE agents.
@@ 5.2 The Costs of Summarization
A closer examination of the cost breakdown reveals that the efficiency gap between strategies stems from two complementary effects. First, the “trajectory elongation” effect discussed in Section˜4.4, and second, the costs of generating the summary.
As shown in Table˜2, the direct API cost of generating summaries accounts for up to 7.2% of the total instance cost. Importantly, these summarization calls are particularly expensive because each requires processing a unique sequence of turns, limiting cache reuse to the LLM-Summary system prompt (Figure˜11). This poses a critical limitation given that, several modern LLM APIs (e.g., Gemini) offer substantially cheaper cache hits than cache misses (up to $10\times$ cheaper). Once we subtract these summarization costs from the total, the efficiency difference between LLM-Summary and Observation Masking largely disappears for most experiments. This indicates that the summarization API calls themselves constitute a substantial portion of the efficiency gap. Nonetheless, the more complex LLM-Summary strategy is still unable to significantly or consistently outperform the remarkably strong Observation Masking strategy on cost-efficiency. This hints at potentially untapped cost savings through underexplored Observation Masking or hybrid approaches.
@@
5.3 Hybrid: Combined Observation Masking and LLM-Summary
Motivated by the strong individual performances of the context management strategies we cover, both significantly and consistently reducing cost by $50\%$, we present a novel hybrid approach in this section. For this, we experiment with the strongest model with respect to solve rate we cover, Qwen3-Coder 480B with SWE-agent [35] on SWE-bench Verified-50 [3] due to cost reasons. We visualize the results in Figure˜5b.
Figure 5: (a) Probing the generality of our findings with OpenHands [30] on the SWE-bench Verified-50 [3] subset. After appropriately tuning the rolling window size $M$ to the agent scaffold, Observation Masking again matches the performance of LLM-Summary on both cost and solve rate. (b) Our novel hybrid Observation Masking and LLM-Summary approach on SWE-bench Verified-50 [3] on the SWE-agent scaffold using Qwen3-Coder 480B. Effectively combining the strengths of each approach results in a strategy that robustly realizes efficiency gains regardless of trajectory length while benefiting from bounded context on excessively long trajectories. Our hybrid approach yields a slight solve-rate gain of 2.6 percent points compared to the Raw Agent while reducing costs by 7% and 11% compared to Observation Masking and LLM-Summary, respectively.
Both the trajectory elongation effect (Section˜4.4) and the summary generation overhead (Section˜5.2) motivate us to treat LLM-Summary as a last resort strategy and defer it as long as possible. However, if we increase $N$, we increase the number of warm-up turns needed until we observe any effect. During this accumulation phase, we operate under the costly Raw Agent regime. Moreover, we risk not observing any effects on short to medium length trajectories (Appendix˜B). Using Observation Masking during the turn accumulation phase allows us to combine the strengths of each approach. By increasing $N$, we defer LLM-Summary and treat this approach as a last resort for bounding the context of long trajectories. At the same time, Observation Masking quickly realizes gains during turn accumulation. Moreover, it does so robustly even on short trajectories.
We set $N=43$, because at this number of turns the context accumulated under the Observation Masking regime approximately matches the context accumulated under the Raw Agent at $N=21$ turns ($\approx 30K$ tokens, see Figure˜9). To avoid notation clash, we use $W$ for the rolling window size of Observation Masking in the hybrid setup. Overall, we use $N=43,M=W=10$ for the hybrid setup. Note that we pass the unmasked context whenever summarizing with LLM-Summary.
Compared to Observation Masking and LLM-Summary, this approach reduces costs by 7% and 11%, respectively. Moreover, it even improves the downstream task performance by 2.6 percent points, pushing the effectiveness-efficiency frontier. This results in expected savings of \$20 compared to Observation Masking and \$35 compared to LLM-Summary on the full SWE-bench Verified [5] benchmark.
To ablate our hyperparameter choice, we also test the hybrid approach with a naive choice of hyperparameters, disregarding any strategy-specific properties. For this, we simply re-use the hyperparameters from the individual approaches $N=21,M=W=10$. With this configuration, the hybrid approach actually degrades the systems’ overall cost efficiency, due to compounding KV cache inefficiencies, and cost overhead due to LLM-Summary invocations.
@@ 6 Limitations
While our study provides a rigorous evaluation of context management strategies, its scope has three main limitations. First, we experiment exclusively within the SE domain, using the SWE-bench [11] benchmark. This domain is characterized by long, verbose tool outputs, a condition that naturally favors the efficiency of Observation Masking. Consequently, our findings on the superiority of this strategy may not generalize to domains where agent-environment interactions are more succinct. Second, all strategies investigated use simple, non-adaptive heuristic triggers. Observation Masking employs a fixed-size rolling window that is agnostic to the relevance or staleness of past observations (e.g., retaining a file’s content after it was modified). Similarly, LLM-Summary operates on a fixed turn-based schedule, ignoring semantic boundaries or agent subgoals. Finally, while we provide initial evidence of generalization across agent scaffolds, a more comprehensive investigation may be warranted.
@@ 7 Conclusion
This work presents a comprehensive study on context management strategies spanning diverse model configurations and agent scaffolds. We find that efficient context management strategies consistently and significantly reduce system costs by $50\%$ without significantly reducing downstream performance. Surprisingly, the popular LLM-Summary strategy is unable to consistently or significantly outperform the simple Observation Masking baseline. This hints at untapped savings potential in modern agentic systems that focus only on LLM-Summary. We empirically validate this hypothesis with our novel hybrid Observation Masking and LLM-Summary strategy that further reduces costs by 7% and 11% compared to Observation Masking and LLM-Summary while improving the downstream task performance. These findings establish the critical need for context management to enable economically feasible, and environmentally sustainable LLM agent deployment. Moreover, it highlights that in the quest for efficient LLM agents, simple solutions can be surprisingly effective.
Acknowledgments and Disclosure of Funding
We would like to thank Calvin Smith for the productive and encouraging discussions on the context management strategies implemented in OpenHands. We would also like to thank Kirill Gelvan for his feedback on early versions of this paper.
@@ Appendix A Experimental Configuration
To compute the instance cost reported in our experiments, we distinguish between proprietary models we access through the API, and models we host locally on our infrastructure. For the Gemini [6] experiments, which we access through the Vertex AI API, we report the cost returned by the API. For Qwen [27, 8] experiments, we self-host the models as described in Section˜3.2. To compute the cost, we use the observed per-turn token usage and post-hoc compute the cost based on the official Alibaba API pricing [^6] https://www.alibabacloud.com/help/en/model-studio/models#16ff9753e1ctz. Note that in contrast to the Vertex AI platform, the official Alibaba API pricing for the Qwen3-32B model, does not distinguish between cache hit and miss input tokens. This leads to an inflated cost of the Qwen3-32B experiments. We refer readers interested in the cost structure of our experiments under different pricing schemes to out HuggingFace dataset [^7] Data: https://huggingface.co/datasets/JetBrains-Research/the-complexity-trap, in which we release all experimental data for our main experiments and our repository [^8] Code: https://github.com/JetBrains-Research/the-complexity-trap.
Appendix B Short Trajectories in Qwen3-32B (thinking): Hypotheses and Implications
The Qwen3-32B (thinking) configuration exhibits an around 50% shorter median trajectory length compared to the Qwen3-32B configuration. Given that the only change between the two configurations is that we enable model thinking, this is surprising. In the following, we detail our investigation of this outlier result and discuss its implications on the interpretation of our findings.
Configuration validation.
We verified that the configuration and dataset are correct for these two experimental configurations across all context management strategies. Additionally, we did not observe deviations in the distribution of instance exit statuses between the two configurations, besides the slightly improved performance of the model under the thinking regime (LABEL:chapter:results).
Qualitative analysis.
As we did not identify a misconfiguration, we further performed a qualitative analysis of the same 20 (4%) trajectories across the Qwen3-32B experimental configurations. Here, we noticed that both models sometimes struggle with following the function calling format of the agent scaffold. However, as discussed above, we did not observe suspicious deviations in the number of exits due to function calling errors in either configuration.
Implications.
Because we could not identify any misconfiguration or otherwise suspicious behavior as the reason for this outlier result with the Qwen3-32B (thinking) configuration, we must assume that it is valid. Thus, we now discuss the implications of this result on the interpretation of our overall findings. First, recall that this configuration resulted in short trajectories. However, both of our context management strategies need a number of warm-up turns, before they start modifying the trajectory and thus reducing cost. For LLM-Summary we require $N+M=31$ turns before we produce the first summary with our hyperparameters. This means the trajectories with a median length of 15 turns are far too short to realize efficiency gains from this context management strategy. Observation Masking on the other hand, starts masking observations after $M=10$ turns. This means we expect to see an effect when using this strategy on shorter trajectories, however it may be muted. This exactly matches the empirical behavior observed in LABEL:chapter:results. Therefore, we attribute the insignificant cost savings in the Qwen3-32B (thinking) configuration to the shorter trajectory lengths, rather than fundamental issues with our context management strategies. This interpretation is further supported by the fact that even in this unfavorable setting, Observation Masking reduces cost by $\approx 10\%$, and both strategies still result in stable downstream performance.
@@ Appendix C Detailed Main Results
Table 3: Comparison of context management strategies with 95% bootstrap confidence intervals, showing asymmetry. We use † to indicate significance compared to the Raw Agent. We report Solve Rate (effectiveness, $\uparrow$) and Instance Cost (efficiency, $\downarrow$). For each model, we boldface the best-performing context management strategy for each metric (relative to the Raw Agent baseline). Change is reported relative to the Raw Agent baseline. All experiments use SWE-agent [35] on SWE-bench Verified [5].
@@ Appendix D Additional Studies
For the critic-enhanced summarizer in Figure˜7, we experiment on SWE-bench Lite-50 [3]. For the sensitivity to the rolling window size $M$ of the Observation Masking strategy and the configurations of the LLM-Summary strategy, we show our results on a randomly sampled 150-instance subset of SWE-bench Verified [5] that we release with our code. We conduct these studies with GPT-4.1-mini [21].
Figure 6: Downstream task performance of a single experiment on a randomly generated 150-sample subset of SWE-bench Verified [5] across various different configuration combinations with respect to the tail length $M$. We find that a larger summarization window compared to the tail length improves performance.
D.1 Observation Masking Configuration
We experiment with the rolling window size $M$ of the Observation Masking strategy. In Figure˜10 we can see that the performance of the strategy peaks at $M=10$, before falling again when further increasing the window size to $M=20$. Thus, we use this configuration of the Observation Masking strategy for our main experiments.